OpenStudy (agentnao):

A ball is dropped from the top of a 1,000 foot building. The height of the ball is half its original height after each bounce. What will be the height of the ball after 9 bounces.

OpenStudy (agentnao):

It's a geometric series and I have the formula to solve it, but I'm not getting the right answer.

OpenStudy (lalaly):

1000, 500, 250 ,125 500/100 =1/2, 250/500=1/2 ar^(n-1)=1000(1/2)^(9-1)=3.90625 feet

OpenStudy (anonymous):

\[1000(1/2)^n\]

OpenStudy (anonymous):

It is geometric sequence

OpenStudy (agentnao):

*sequence, yes. Sorry.

OpenStudy (lalaly):

suposed to be 500/1000 second line